# Recurrence Criteria for Reducible Homogeneous Open Quantum Walks on the Line

### v2

@@ Abstract
In this paper, we study the recurrence of homogeneous Open Quantum Walks (HOQWs) induced by finite-dimensional coins $(L,B,R)$. We present three distinct recurrence criteria, each adapted to different types of coins. The first criterion applies to coins whose auxiliary map has a unique invariant state, resulting in the first recurrence criterion for Lazy HOQWs. The second one applies to Lazy HOQWs of dimension 2, where we provide a complete characterization of the recurrence for this low-dimensional case. Finally, we present a general criterion for finite-dimensional coins in the non-lazy case $(B=0)$, which generalizes many of the previously known results. This new criterion holds for irreducible and reducible HOQWs.
@@ 1 Introduction
Open Quantum Walks (OQWs) naturally extend classical Markov chains within the framework of open quantum systems. Originally proposed by [2], OQWs describe open quantum systems in which the repeated measurement of the particle’s position, together with the space of particle states, gives rise to a Markov chain.
In this context, we consider homogeneous OQWs (HOQWs) defined on the set of integers, that is, the graph of this type of random walk is the set of vertices $V=\mathbb{Z}$, where 3 linear operators $L,B,R$ play the roles of transitions between the vertices of the quantum walk, which are constant and each jump only occurs from one vertex to its nearest neighbors, or to itself. HOQWs on more general lattices are explored and detailed in [7].
Recurrence is an important probabilistic property of any random walk, often characterized by the expected number of returns to the initial vertex, emerging as a central question of interest in the quantum context, and can be directly connected with the classical recurrence criteria for random walks in $\mathbb{Z}$. For example, consider the classical homogeneous random walk on $\mathbb{Z}$ with graph given by Figure 1. It is well known that this lazy random walk with $r,l\geq 0,\;b>0,\;l+b+r=1,$ is recurrent if and only if the transition parameters $r$ (rate of jumping to right) and $l$ (rate of jumping to left) are equal. Also, if $b=0,$ then the walk is non-lazy and it is recurrent if and only if $r=1/2=l$ [14].
@@
With this description, it is natural to seek criteria to determine under what conditions a HOQW is recurrent given $L,B,R.$ This problem has been addressed in the literature with the development of specific criteria, which are essential to understanding the dynamics and long-term behavior of these systems. See for instance [8, 13, 12]. These works define recurrence in terms of probability of return, where a state is said to be recurrent if the probability of the walk returning to it is equal to 1. However, it has been shown that for OQWs, the two definitions of recurrence based on returning with probability 1 is not equivalent to the one based on the average number of returns being infinite [3]. This property highlights a distinct characteristic of OQWs compared to classical Markov chains.
Once this difference is understood, we discuss the recurrence criteria for OQWs, adopting the definition proposed in [3]. Specifically, we analyze the recurrence of OQWs based on the average number of returns.
@@
A first Central Limit Theorem (CLT) for Open Quantum Walks (OQWs) was established in [1]. In this work, we rely on more recent developments, including the CLT for Lazy OQWs presented in [10], as well as the general CLT for homogeneous OQWs (HOQWs) on lattices with finite internal degrees of freedom given in [6].
This work is structured as follows. In Section 2 we define the open quantum walks induced by a coin and their transition probabilities. Section 3 is devoted to presenting a recurrence criterion for a class of Lazy HOQWs, and one restricted to Lazy HOQWs of dimension 2. In Section 4 we focus on the case where the coin is non-lazy and we present a general recurrence criterion for finite-dimensional coins. Section 5 illustrates the results with comprehensive examples.
@@ 2 General Settings
If $B\neq 0,$ the quantum walk exhibits a lazy behavior, meaning that in the context of quantum walks there is a non-zero probability for the walker to stay in the same vertex after a step. We refer to the HOQW above as a HOQW induced by the coin $(L,B,R)$ in which the walk is inherently lazy.
Therefore, this coin represents a homogeneous, nearest-neighbor OQW on the integer line with the left and right transitions given by $L$ and $R$, respectively, and $B$ represents the rate of no jump at some instant.
This lazy behavior introduces distinct properties compared to its non-lazy counterpart. Notably, the presence of the matrix $B$ influences the walk’s propensity to remain in its current vertex, thereby affecting the overall dynamics. If $B=0,$ then $\Phi$ will be called a HOQW induced by a coin $(L,R)$ or a non-lazy HOQW induced by a coin $(L,R).$ The dynamic described above is represented by the graph in Figure 2.
@@
The quantum trajectory of this kind of OQW on the line, induced by a coin $(L,B,R)$ and starting from a state $\tau$ of the form $\tau=\sum_{i\in\mathbb{Z}}\tau_{i}\otimes|i\rangle\langle i|,$ is any path generated by the Markov chain $(X_{n},\tau_{n})_{n\geq 0}$, where $X_{n}$ denotes the position of the particle at time $n$ and $\tau_{n}$ its internal degree. The transition probabilities are given by
@@ 3.1 Faithfulness of Density Operators
Let us consider an OQW starting the walk on vertex $\lvert i\rangle$ with initial density operator $\rho.$ For evolution on time $n=0,1,2,\ldots,$ we define
$p_{n}(j)=\mathbb{P}(\mbox{the quantum particle, at time }n,\mbox{ is in site}\;\lvert j\rangle)=\mathrm{Tr}(\rho_{n}(j)),$
@@
$p_{ji;\rho}(n)=p_{n}(\rho\otimes|i\rangle\to|j\rangle)=\mathrm{\mathrm{Tr}}(\rho_{n}(j)\otimes\lvert j\rangle\langle j\rvert)=\mathrm{\mathrm{Tr}}\left(\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert)(I\otimes\lvert j\rangle\langle j\rvert)\right)$
@@
Consider an OQW on some set of vertices $\mathbb{Z}$, let $i\in\mathbb{Z}$ and $\rho\in\mathcal{D}(\mathfrak{h})$. We say that a vertex $\lvert i\rangle$ is
@@
    transient, if $\lvert i\rangle$ is $\rho-$transient for all $\rho\in\rho\in\mathcal{D}(\mathfrak{h}).$
@@
$\infty=\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\tau\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]=\sum_{j}\left(\lambda_{j}\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\lvert e_{j}\rangle\langle e_{j}\rvert\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]\right),$
@@
Let $\Phi$ be a HOQW on $\mathbb{Z}$ with $\dim(\mathfrak{h})=d<\infty.$ If $\Phi$ is $\tau$-recurrent for some $\tau\in\mathcal{D}(\mathfrak{h}),$ then
@@ 3.2 Enclosures and Invariant States
A minimal enclosure with finite dimension is the support of a unique invariant state of the channel, that is, a density operator $\sigma$ such that $\mathcal{L}(\sigma)=\sigma$. For more on enclosures of OQWs, we refer the reader to [6, 7, 3].
@@
Consider a HOQW induced by a coin $(L,B,R)$ of dimension $d$. We are in one (and only one) of the following situations:
@@
Moreover, by equation (4) we have that $\cup_{n}\;\text{supp}(\rho_{n})$ is a subspace of $\mathfrak{h}.$ Joining all the possible initial states in which $(L,B,R)$ is transient, we obtain a decomposition (3), where $\mathcal{Y}$ or $\mathcal{X}$ may be null, and by definition $\mathcal{Y}$ is an enclosure.
@@
This result can be obtained by using a Doob decomposition as described in [10, Equation 27] and [1, Theorem 5.2] for a Martingale $M_{n}$ satisfying
@@
Consider a HOQW induced by a coin $(L,B,R),\;\mathcal{L}$ its auxiliary map and $\mathcal{Y}$ an enclosure. Suppose that the initial density operator is $\rho\in\mathcal{D}(\mathcal{Y})$ with quantum trajectories $(X_{n},\rho_{n})_{n\geq 0}$ and $X_{n}/n\rightarrow 0$ in probability, then $(L,B,R)$ is $\sigma$-recurrent for some $\sigma\in\mathcal{D}(\mathcal{Y}).$
@@
The proof is analogous to [11, Theorem 7] and we will proceed by remarking on how it can be generalized. Firstly, the maximum appearing in [11, Lemma 6] can be taken in $\mathcal{D}=\mathcal{D}(\mathcal{Y}).$ Indeed, if the initial state is $\rho\in\mathcal{D}(\mathcal{Y}),$ then $\rho_{n}\in\mathcal{D}(\mathcal{Y})$ for every $n,$ thereby the maximum taken in the proof of [11, Lemma 6] can be reduced to $\mathcal{D}(\mathcal{Y}).$ Therefore, by the same reason, the density operator considered in [11, Theorem 7] belongs to $\mathcal{D}(\mathcal{Y}),$ proving the Theorem for the non-lazy case. Moreover, the lazy case holds as an extension of this proof, since all the results work for homogeneous OQWs on $\mathbb{Z}$.
@@
The following theorem presents the first recurrence criterion for non-irreducible HOQWs of arbitrary finite dimension. This result extends [11, Corollary 8 and Proposition 8] for reducible CTOQWs and can also be applied to Lazy HOQWs.
@@
Let us consider a HOQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension $d$ such that the auxiliary map $\mathcal{L}$ is ergodic. Then $\mathcal{L}$ has a unique invariant state $\rho_{\infty}$ and
@@
If $\;\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then $X_{n}/n\rightarrow 0$ by Theorem 3.5, thus $(L,B,R)$ is $\rho$-recurrent for some $\rho\in\mathcal{D}(\mathcal{Y})$ by Theorem 3.6, which is a contradiction.
[…]
@@ 3.3 General Recurrence Criteria for HOQWs induced by a coin $(L,B,R)$ of dimension 2
Building upon the established preceding proposition, we will advance to the following theorem, delineating a comprehensive criterion for 2-dimensional coins, extending [11, Theorem 17] to encompass the Lazy case. This criterion will be structured into two distinct parts: the first adhering to the logic of the previous theorem, and the second articulated in terms of the eigenvalues of the operators of the coin, independent of the invariant densities of the auxiliary map.
@@
Consider a HOQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension 2.
@@
$(2.1)\quad|l_{1}|=|r_{1}|\mbox{ and }|l_{2}|=|r_{2}|$$\Rightarrow$ $(L,B,R)\mbox{ is recurrent,}$,
$(2.2)\quad|l_{1}|\neq|r_{1}|\mbox{ and }|l_{2}|\neq|r_{2}|$$\Rightarrow$ $(L,B,R)\mbox{ is transient,}$,
$(2.3)\quad|l_{i}|=|r_{i}|\mbox{ and }|l_{j}|\neq|r_{j}|$$\Rightarrow$ $(L,B,R)\mbox{ is transient with respect to }\sigma_{j}=\lvert u_{j}\rangle\langle u_{j}\rvert\mbox{ and it is recurrent}$
@@
Analogously, if $|l_{1}|\neq|r_{1}|$ and $|l_{2}|\neq|r_{2}|,$ then for $\rho_{1}=\lvert u_{1}\rangle\langle u_{1}\rvert,$ and $\rho_{2}=\lvert u_{2}\rangle\langle u_{2}\rvert,$ where $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ are orthonormal common eigenvectors of $L,B,R,$ we have $p_{j,j+1;\rho}(n)<p_{j,j-1;\rho}(n)$ (or $>$) for all $j\in\mathbb{Z}$ and all $n\geq 0$ and therefore in this case we have a classical homogeneous walk which is transient for both $\rho_{1},\rho_{2}$. Since the walk is transient for two distinct eigenvectors, the walk is transient by corollary 3.2.

### v3

@@ Abstract
In this paper, we study the recurrence of Open Quantum Walks induced by finite-dimensional coins on the line ($\mathbb{Z}$) and on the grid ($\mathbb{Z}^{2}$). Two versions are considered: discrete-time open quantum walks (OQW) and continuous-time open quantum walks (CTOQW). We present three distinct recurrence criteria for OQWs on $\mathbb{Z}$, each adapted to different types of coins. The first criterion applies to coins whose auxiliary map has a unique invariant state, resulting in the first recurrence criterion for Lazy OQWs. The second one applies to Lazy OQWs of dimension 2, where we provide a complete characterization of the recurrence for this low-dimensional case. The third one presents a general criterion for finite-dimensional coins in the non-lazy case, which generalizes many of the previously known results for OQWs on $\mathbb{Z}$. Also, we present a general recurrence criterion for OQWs on $\mathbb{Z}^{2}$ via the open quantum jump chain, obtained from a recurrence criterion for CTOQWs on $\mathbb{Z}^{2}$.
@@ 1 Introduction
Open Quantum Walks (OQWs) naturally extend classical Markov chains within the framework of open quantum systems. Originally proposed by [2], OQWs describe open quantum systems in which repeated measurements of the particle’s position, together with the particle’s state space, give rise to a Markov chain.
In this context, we first consider homogeneous Open Quantum Walks (HOQWs) defined on the set of integers, that is, the graph of this type of walk has the set of sites $V=\mathbb{Z}$, where three linear operators $L,B,R$ correspond to transitions between sites in the quantum walk, remaining constant, with each step moving only to the nearest neighbors $(R,L)$ or staying at the same site $(B)$.
Recurrence is an important probabilistic property of any walk, often characterized by the expected number of returns to the initial site, bringing a central concern in the quantum context, and can be directly related to classical recurrence criteria for walks in $\mathbb{Z}$. For example, consider the classical homogeneous walk on $\mathbb{Z}$ with graph as shown in Figure 1.
It is well known that this lazy random walk with $r,l\geq 0,\;b>0,\;l+b+r=1,$ is recurrent if and only if the transition parameters $r$ (rate of jumping to the right) and $l$ (rate of jumping to the left) are equal. Also, if $b=0,$ then the walk is non-lazy and it is recurrent if and only if $r=1/2=l$ [21].
@@
With this description, it is natural to seek criteria to determine under what conditions an HOQW is recurrent, given $L,B,R.$ This problem has been addressed in the literature with the development of specific criteria, which are essential to understanding the dynamics and long-term behavior of these systems. See, for instance, [12, 17, 16]. These works define recurrence in terms of probability of return, where a state is said to be recurrent if the probability of the walk returning to it is equal to 1. However, it has been shown that for OQWs, the two definitions of recurrence based on returning with probability 1 are not equivalent to the one based on the average number of returns being infinite [4]. This property highlights a distinct characteristic of OQWs compared to classical Markov chains.
Once this difference is understood, we discuss the recurrence criteria for HOQWs, adopting the definition proposed in [4]. Specifically, we analyze the recurrence of HOQWs based on the average number of returns.
@@
A first Central Limit Theorem (CLT) for Open Quantum Walks (OQWs) was established in [1]. In this work, we rely on more recent developments, including the CLT for Lazy HOQWs presented in [14], as well as the general CLT for HOQWs on lattices with finite internal degrees of freedom given in [10].
By making use of the ideas above and generalizing results from the classical version to recurrence, we also obtain a recurrence criterion for the HOQW on the grid, which is the case where the set of vertices is $\mathbb{Z}^{2}$; however, this is achieved through continuous-time OQWs (CTOQWs). For this purpose, we use recent results by [20], where the introduction of the open quantum jump chain allowed for the association between recurrence in discrete-time and continuous-time versions of OQWs.
This work is structured as follows. In Section 2, we define the open quantum walks induced by a coin and their transition probabilities. Section 3 is devoted to presenting a recurrence criterion for a class of Lazy HOQWs, and one restricted to Lazy HOQWs of dimension 2. In Section 4, we focus on the case where the coin is non-lazy, and we present a general recurrence criterion for finite-dimensional coins. Section 5 illustrates a first pack of comprehensive examples. Section 6 discusses OQWs on the grid, and Section 7 brings examples of HOQWs on those walks.
@@ 2 General Settings
If $B\neq 0,$ the quantum walk exhibits a lazy behavior, meaning that in the context of quantum walks, there is a non-zero probability for the walker to stay in the same vertex after a step. We refer to the HOQW above as a HOQW induced by the coin $(L,B,R)$ in which the walk is inherently lazy.
Therefore, this coin represents a homogeneous, nearest-neighbor OQW on the integer line with the left and right transitions given by $L$ and $R$, respectively, while $B$ describes the rate of no jump at some instant.
This lazy behavior introduces distinct properties compared to its non-lazy counterpart. Notably, the presence of the matrix $B$ influences the walk’s propensity to remain in its current vertex, thereby affecting the overall dynamics. If $B=0,$ then $\Phi$ will be called a HOQW induced by a coin $(L,R),$ or a non-lazy HOQW induced by a coin $(L,R).$ The dynamic described above is represented by the graph in Figure 2.
@@
The quantum trajectory of this kind of HOQW on the line, induced by a coin $(L,B,R)$ and starting from a state $\tau$ of the form $\tau=\sum_{i\in\mathbb{Z}}\tau_{i}\otimes|i\rangle\langle i|,$ is any path generated by the Markov chain $(X_{n},\tau_{n})_{n\geq 0}$, where $X_{n}$ denotes the position of the particle at time $n$ and $\tau_{n}$ its internal degree. The transition probabilities are given by
@@ 3.1 Faithfulness of Density Operators
Let us consider an HOQW starting the walk on vertex $\lvert i\rangle$ with initial density operator $\rho.$ For evolution on time $n=0,1,2,\ldots,$ we define
[…]
@@
[…]
@@
Consider an HOQW on some set of vertices $\mathbb{Z}$, let $i\in\mathbb{Z}$ and $\rho\in\mathcal{D}(\mathfrak{h})$. We say that a vertex $\lvert i\rangle$ is
@@
    transient, if $\lvert i\rangle$ is $\rho$-transient for all $\rho\in\mathcal{D}(\mathfrak{h}).$
@@
[…]
@@
Let $\Phi$ be an HOQW on $\mathbb{Z}$ with $\dim(\mathfrak{h})=d<\infty.$ If $\Phi$ is $\tau$-recurrent for some $\tau\in\mathcal{D}(\mathfrak{h}),$ then
@@ 3.2 Enclosures and Invariant States
A minimal enclosure with finite dimension is the support of a unique invariant state of the channel, that is, a density operator $\sigma$ such that $\mathcal{L}(\sigma)=\sigma$. For more on enclosures of HOQWs, we refer the reader to [10, 11, 4].
@@
Remark 3.4.
From now on we disregard the trivial cases where at least one of the operators $L,B,R$ has some eigenvalue $\lambda$ of modulus 1. To clarify, let $T\lvert v\rangle=\lambda\lvert v\rangle,$ where $\lvert v\rangle$ is an unitary vector and $T\in\{L,B,R\}.$ In this case, starting the walk with density operator $\rho=\lvert v\rangle\langle v\rvert$, the walker is either absorbed by a vertex $(T=B)$ or the particle always jumps to the same side $(T=L$ or $T=R).$
@@
Consider an HOQW induced by a coin $(L,B,R)$ of dimension $d$. We are in one (and only one) of the following situations:
@@
Moreover, by equation (4) we have that $\cup_{n}\;\text{supp}(\rho_{n})$ is a subspace of $\mathfrak{h}.$ Joining all the possible initial states in which $(L,B,R)$ is transient, we obtain a decomposition (3), where $\mathcal{Y}$ or $\mathcal{X}$ may be $\{0\}$, and by definition $\mathcal{Y}$ is an enclosure.
@@
This result can be obtained by using a Doob decomposition as described in [14, Equation 27] and [1, Theorem 5.2] for a Martingale $M_{n}$ satisfying
@@
Consider an HOQW induced by a coin $(L,B,R),\;\mathcal{L}$ its auxiliary map and $\mathcal{Y}$ an enclosure. Suppose that the initial density operator is $\rho\in\mathcal{D}(\mathcal{Y})$ with quantum trajectories $(X_{n},\rho_{n})_{n\geq 0}$ and $X_{n}/n\rightarrow 0$ in probability, then $(L,B,R)$ is $\sigma$-recurrent for some $\sigma\in\mathcal{D}(\mathcal{Y}).$
@@
The proof is analogous to [15, Theorem 7] and we will proceed by remarking on how it can be generalized. Firstly, the maximum appearing in [15, Lemma 6] can be taken in $\mathcal{D}=\mathcal{D}(\mathcal{Y}).$ Indeed, if the initial state is $\rho\in\mathcal{D}(\mathcal{Y}),$ then $\rho_{n}\in\mathcal{D}(\mathcal{Y})$ for every $n,$ thereby the maximum taken in the proof of [15, Lemma 6] can be reduced to $\mathcal{D}(\mathcal{Y}).$ Therefore, by the same reason, the density operator considered in [15, Theorem 7] belongs to $\mathcal{D}(\mathcal{Y}),$ proving the Theorem for the non-lazy case. Moreover, the lazy case holds as an extension of this proof, since all the results work for HOQWs on $\mathbb{Z}$.
@@
The following theorem presents the first recurrence criterion for non-irreducible HOQWs of arbitrary finite dimension. This result extends [15, Corollary 8 and Proposition 8] for reducible CTOQWs and can also be applied to Lazy HOQWs.
@@
Let us consider an HOQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension $d$ such that the auxiliary map $\mathcal{L}$ is ergodic. Then $\mathcal{L}$ has a unique invariant state $\rho_{\infty}$ and
@@
If $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right)$ then $X_{n}/n\rightarrow 0$ by Theorem 3.6, thus $(L,B,R)$ is $\rho$-recurrent for some $\rho\in\mathcal{D}(\mathcal{Y})$ by Theorem 3.7, which is a contradiction.
On the other hand, if $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)\neq\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then $X_{n}/n\rightarrow m\neq 0.$ Therefore $\mathbb{P}_{0,\rho}(X_{n}=0\mbox{ i.o.})=0,$ showing that the mean number of returns to the initial site is finite for any $\rho$, thus the walk is transient, which is also a contradiction. We have shown that either the walk is recurrent or transient.
@@ 3.3 General Recurrence Criteria for HOQWs induced by a coin $(L,B,R)$ of dimension 2
Building upon the established preceding proposition, we will advance to the following theorem, delineating a comprehensive criterion for 2-dimensional coins, extending [15, Theorem 17] to encompass the Lazy case. This criterion will be structured into two distinct parts: the first adhering to the logic of the previous theorem, and the second articulated in terms of the eigenvalues of the operators of the coin, independent of the invariant densities of the auxiliary map.
@@
Consider an HOQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension 2.
@@
$(2.1)\quad|l_{1}|=|r_{1}|\mbox{ and }|l_{2}|=|r_{2}|$ $\Rightarrow$ $(L,B,R)\mbox{ is recurrent,}$
$(2.2)\quad|l_{1}|\neq|r_{1}|\mbox{ and }|l_{2}|\neq|r_{2}|$ $\Rightarrow$ $(L,B,R)\mbox{ is transient,}$
[…]
@@
Analogously, if $|l_{1}|\neq|r_{1}|$ and $|l_{2}|\neq|r_{2}|,$ then for $\rho_{1}=\lvert u_{1}\rangle\langle u_{1}\rvert$, and $\rho_{2}=\lvert u_{2}\rangle\langle u_{2}\rvert$, where $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ are orthonormal common eigenvectors of $L,B,R,$ we have $p_{j,j+1;\rho}(n)<p_{j,j-1;\rho}(n)$ (or >>) for all $j\in\mathbb{Z}$ and all $n\geq 0$ and therefore in this case we have a classical homogeneous walk which is transient for both $\rho_{1},\rho_{2}$. Since the walk is transient for two distinct eigenvectors, the walk is transient by corollary 3.2.